# Beam tracing and profile evolution for localised beams in inhomogeneous plasmas

### v2

@@ IV Dissipation and heating
where $\bm{\epsilon}_{\mathrm{H}}$ is the Hermitian part of the plasma dielectric tensor and $\bm{\epsilon}_{\mathrm{A}}$ is some anti-Hermitian contribution responsible for heating of resonant particlesvan1993dielectric. Note that other approaches exist for modelling non-Hermitian dielectric tensorsFriedland1980. For simplicity, we assume that $|\bm{\epsilon}_{\mathrm{A}}|\sim\frac{\lambda}{L}|\bm{\epsilon}_{\mathrm{H}}|$. This is to ensure that $\bm{\epsilon}_{\mathrm{A}}$ only enters at second order, such that the Hamiltonian is strictly real to lowest order. With this ordering, hot plasma dissipation appears only as a modification to the profile evolution equation.
$\frac{\partial\mathcal{P}}{\partial\tau}-\mathrm{i}\left(\frac{1}{2}\partial K_{\mu}\partial K_{\nu}H\partial_{\mu}\partial_{\nu}(\mathcal{P})^{(0)}\right)+w_{\mu}T_{\mu\nu}\partial_{\nu}\mathcal{P}^{(0)}+\mathrm{i}\hat{e}^{*}_{\mu}\epsilon_{A,\mu\nu}(\tau,\bm{w})\hat{e}_{\nu}\mathcal{P}=0.$ (81)
$\hat{\bm{e}}$ is the polarisation we obtained from solving the cold-plasma dispersion relation. Notice that due to the small length scale of $\bm{\epsilon}_{\mathrm{A}}$, it is not only a function of $\tau$, but one of $\bm{w}$ as well. This additional width dependence is crucial for calculating power deposition accuratelyPARADE2021. Nonetheless, if it were constant across the width we can easily solve for the amplitude decay, since $\hat{e}^{*}_{\mu}\epsilon_{A,\mu\nu}(\tau,\bm{w})\hat{e}_{\nu}$ is strictly imaginary. Our amplitude will simply be scaled by $\exp\left(-\int_{0}^{\tau}\mathrm{i}\hat{e}^{*}_{\mu}\epsilon_{A,\mu\nu}(\tau^{\prime})\hat{e}_{\nu}\mathrm{d}\tau^{\prime}\right)$, which results in attenuation of our beam due to absorption of energy by electron heating. However, when it depends on $\bm{w}$, we are no longer able to solve this exactly. Thus, numerically solving eq. 81 would still be the only way to approach this.
V Discussion
V.1 Propagation in homogeneous, isotropic media
A useful testbed for eq. 23 and eq. 44 is the case of Gauss-Hermite beams in homogeneous media. We will focus on the results in the vacuum case, as propagation in any homogeneous isotropic medium is virtually identical to propagation through vacuum, just with a different refractive index. Thus, in vacuum(or any homogeneous medium), we will use a positively propagating Hamiltonian, $H=\frac{K^{2}}{K_{0}^{2}}-1$. This is an important sanity check as vacuum-propagation of Gauss-Hermite beams is a well-established textbook result. Various bases of solutions exist in vacuum, the particular bases of solutions we will focus on are the ‘elegant’ Gauss-Hermite beams, given by:
@@ V.1 Propagation in homogeneous, isotropic media
V.1.1 ‘Elegant’ Hermite polynomial basis
@@
and it acts on $\exp\left(\frac{\mathrm{i}}{2}\bm{w}\cdot\bm{\Psi}_{w}\cdot\bm{w}\right)$. This ladder operator once again has the amplitude evolution built right into it. It is now clear that the complex Hermite polynomials are also given by both our ladder operators, and that the evolution of their profiles are identical for all three approaches.
We wish to point out that although the ‘elegant’ Gauss-Hermite beams are biorthogonal under the bilinear pairing inner product(section. III.5.3), this means that the elegant Gauss-Hermite beams do not neatly diagonalise in power, thus it is challenging to analyse the total power across the modes using this basis. An alternative basis of solutions that neatly diagonalises in power is the traditional Gauss-Hermite beam ansatzpampaloni2004.
V.1.2 Traditional Hermite polynomial basis
@@
V.2 Relation to prior work by Pereversev
To the best our knowledge, prior work by PereversevPere1998 is the most comprehensive and directly comparable attempt at beam tracing of an arbitrary profile. He does so by directly substituting in a Hermite polynomial basis of solutions into the Helmholtz equation, then analysing terms of various orders in the resulting expansion. This approach differs from our approach as it immediately prescribes a basis of solutions, whereas our results allow one to use any suitably chosen basis of solutions or even match boundary conditions exactly, simply by tweaking the ladder operators. An important feature to note is that our evolution equations for the beam envelope is decoupled from the profile evolution, whereas it is coupled to the profile evolution in Pereversev’s work. Whilst this may not have the biggest effect numerically, it certainly makes finding analytic solutions much more challenging relative to our approach.
@@ V.2 Relation to prior work by Pereversev
Pereversev does not treat eqs. 19, 20, 18, 23 separately, but rather considers them together by substituting in a specific basis of solutions. Upon substituting in the ansatz:
@@
$\begin{split}&\frac{\mathrm{d}\bm{\Psi}}{\mathrm{d}\tau}+\bm{\Psi}\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}(H)\bm{\Psi}+\bm{\Psi}\bm{\nabla}_{\bm{K}}\bm{\nabla}(H)+\bm{\nabla}\bm{\nabla}_{\bm{K}}(H)\bm{\Psi}\\ &+\bm{\nabla}\bm{\nabla}(H)=\bm{v}_{1}\bm{v}_{1}\cdot\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}(H)\cdot\bm{v}_{1}\bm{v}_{1}+\bm{v}_{2}\bm{v}_{2}\cdot\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}(H)\cdot\bm{v}_{2}\bm{v}_{2},\\ &\frac{\mathrm{dln}(\alpha^{(0)})}{\mathrm{d}\tau}+\frac{\mathrm{d}\bm{K}}{\mathrm{d}\tau}\cdot\bm{\nabla}_{\bm{K}}\hat{e}_{m}\hat{e}^{*}_{m}+\hat{e}^{*}_{i}\partial K_{\mu}D_{im}\partial_{\mu}\hat{e}_{m}+\frac{1}{2}\Psi_{\mu\nu}\partial K_{\mu}\partial K_{\nu}H\\ &+\mathrm{i}(2n+1)\bm{v}_{1}\cdot\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}(H)\cdot\bm{v}_{1}+\mathrm{i}(2m+1)\bm{v}_{2}\cdot\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}(H)\cdot\bm{v}_{2}=0.\end{split}$ (99)
@@
$\frac{\mathrm{d}^{n}}{\mathrm{d}\tau^{n}}(v_{1}^{\mu}\partial K_{\mu}\partial K_{\nu}Hv_{2}^{\nu})|_{\tau=0}=0,$ (100)
implying that $v_{1}^{\mu}\partial K_{\mu}\partial K_{\nu}Hv_{2}^{\nu}=0$ throughout the evolution if it is 0 at the boundary. Firstly, note that this solution is a simple product of Hermite polynomials without arguing that they diagonalise $\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H$, as if they did they would strictly be orthogonal to each other. However, we argued earlier that they have to be orthogonal in order for a simple product of Hermite polynomials to be the solution in our basis. However, it is still entirely possible that a sum of Hermite polynomial products could neatly reduce to a single product. Since Pereversev’s arguments in the Hermite polynomials differ from our arguments, it may still be the case that they are identical. In homogeneous media they are certainly identical.
@@
$v_{1}^{\mu}v_{2}^{\nu}\left(\frac{\mathrm{d}}{\mathrm{d}\tau}(\partial K_{\mu}\partial K_{\nu}H)\right)_{w}=v_{1}^{\mu}v_{2}^{\nu}(T_{w,\mu\rho}\left(\partial K_{\rho}\partial K_{\nu}H\right)_{w}+T_{w,\nu\rho}\left(\partial K_{\rho}\partial K_{\mu}H\right)_{w})$ (101)
@@
Since $\bm{v}_{I}$ ‘diagonalise’ $(\partial K_{\mu}\partial K_{\nu}H)_{w}$ in vacuum, and remain orthogonal to each other throughout their evolution, Pereversev’s solution reproduces the exact same solutions that we obtained earlier in vacuum. This follows naturally from our derivation of Pereversev’s beam tracing equations from our profile evolution equation in eq. 99. Finding a situation in which our solutions do not coincide would prove that eq. 100 is not satisfied in general. We can only do so in a system where $\bm{v}_{I}$ do not diagonalise $(\partial K_{\mu}\partial K_{\nu}H)_{w}$, and one good place to look for this is in propagation where the vectors rotate about $\hat{\bm{g}}$. Only by comparing the predictions of our theory with Pereversev’s in that context would one be able to investigate if our theories do in fact give identical results for the Gauss-Hermite boundary conditions. Although analytic work on obtaining the solutions in slab geometrypoli1999paraxial; Tsironis2006 has been done, the vectors do not rotate about $\hat{\bm{g}}$, nor is $(\partial K_{\mu}\partial K_{\nu}H)_{w}$ non-diagonal in the first place. This means that the solutions they obtained will coincide with those obtained by this method, given the same boundary conditions. Unfortunately, we are unable to construct a physical situation where we are able to solve for analytically whilst rotating the vectors, thus this may be a task for numerical comparison.
@@
V.3 Further Work
Apart from theoretical work, work must go into computational implementation of our theory. The beam-tracing code of choice would ScottyHallChen2022, as our theoretical model will simply reside as an additional module on top of its base beam-tracing capabilities for DBS. This work would hopefully enable Gauss-Hermite DBS analysis in the future, in order to provide more accurate signal analysis for imperfect antenna patterns.
In that same vein, we hope to utilise the results of this theory to develop a simplified theory of DBS, and maybe even cross-polarisation scattering(CPS), but with Gauss-Hermite modes – similar to the approach of previous work in DBS focused on the lowest order Gauss-Hermite beamHallChen2022; HallChen2022b; HallChen2024. We hope that this model will be computationally more efficient than evaluating the full volume integral of the reciprocity theorem, such that it is capable of providing real-time synthetic diagnostic analysis. Above all, we believe such a theory would give a clearer account of various factors that can affect the diagnostic signal, enabling accurate extraction of the turbulence spectrum from the signal, beyond just optimising diagnostic design. In order to do this, figuring out discontinuous boundary conditions would also be crucial.
VI Conclusion
@@
We proceed to obtain a basis of solutions to the profile-evolution equation, for any choice of initial profile. These solutions were obtained using ladder operators, which can be used to set a wide range of polynomial boundary profiles. We focused on constructing ladder operators that generated functions which reduced to the Gauss-Hermite beams at the plasma boundary. After investigating several useful properties of the ladder operators, we found that the biorthogonal system of Hermite polynomials and Hermite functions evolves into a biorthogonal set of solutions in the plasma, which is a useful feature for signal analysis. Finally, we utilised our ladder operators to construct the exact solution to the propagation of a Gauss-Hermite beam through an inhomogeneous cold plasma using eq. 79. Most importantly however, this solution was derived directly from solving eq. 23 in full generality, thus we are confident in the accuracy of the result.
We finally demonstrated our ladder operators in action for propagation of a Gauss-Hermite beam through homogeneous media, before discussing future research direction, in particular, hopefully applying the results of this paper for improving real-time microwave diagnostics analysis. We aim to utilise the useful mathematical properties of our solutions obtained in this paper to analyse general diagnostic signals.

### v3

@@ IV Dissipation and heating
where $\bm{\epsilon}_{\mathrm{H}}$ is the Hermitian part of the plasma dielectric tensor and $\bm{\epsilon}_{\mathrm{A}}$ is some anti-Hermitian contribution responsible for heating of resonant particles[van1993dielectric]. Note that other approaches exist for modelling non-Hermitian dielectric tensors[Friedland1980], but we have not found an approach to integrate such an approach into our model. For simplicity, we assume that $|\bm{\epsilon}_{\mathrm{A}}|\sim\frac{\lambda}{L}|\bm{\epsilon}_{\mathrm{H}}|$. This is to ensure that $\bm{\epsilon}_{\mathrm{A}}$ only enters at second order, such that the Hamiltonian is strictly real to lowest order. With this ordering, hot plasma dissipation appears only as a modification to the profile evolution equation, which given as a PDE. Index notation will be utilised as it makes the derivatives explicit without needing to bother oneself with the orderings, it is given by:
$\frac{\partial\mathcal{P}}{\partial\tau}-\mathrm{i}\left(\frac{1}{2}\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}H\frac{\partial}{\partial w_{\mu}}\frac{\partial}{\partial w_{\nu}}(\mathcal{P})^{(0)}\right)+w_{\mu}T_{\mu\nu}\frac{\partial}{\partial w_{\nu}}\mathcal{P}^{(0)}+\mathrm{i}\hat{e}^{*}_{\mu}\epsilon_{A,\mu\nu}(\tau,\bm{w})\hat{e}_{\nu}\mathcal{P}=0.$ (72)
$\hat{\bm{e}}$ is the polarisation we obtained from solving the cold-plasma dispersion relation. Notice that due to the small length scale of $\bm{\epsilon}_{\mathrm{A}}$, it is not only a function of $\tau$, but one of $\bm{w}$ as well. This additional width dependence is crucial for calculating power deposition accurately[PARADE2021]. Nonetheless, if it were constant across the width we can easily solve for the amplitude decay, since $\hat{e}^{*}_{\mu}\epsilon_{A,\mu\nu}(\tau,\bm{w})\hat{e}_{\nu}$ is strictly imaginary. Our amplitude will simply be scaled by $\exp\left(-\int_{0}^{\tau}\mathrm{i}\hat{e}^{*}_{\mu}\epsilon_{A,\mu\nu}(\tau^{\prime})\hat{e}_{\nu}\mathrm{d}\tau^{\prime}\right)$, which results in attenuation of our beam due to absorption of energy by electron heating. However, when it depends on $\bm{w}$, we are no longer able to solve this exactly. Thus, numerically solving eq. 72 would still be the only way to approach this.
V Propagation in homogeneous, isotropic media
A useful testbed for eq. 24 and eq. 46 is the case of Gauss-Hermite beams in homogeneous media. We will focus on the results in the vacuum case, as propagation in any homogeneous isotropic medium is virtually identical to propagation through vacuum, just with a different refractive index. Thus, in vacuum(or any homogeneous medium), we will use a positively propagating Hamiltonian, $H=\frac{K^{2}}{K_{0}^{2}}-1$. This is an important consistency check as vacuum propagation of Gauss-Hermite beams is a well established textbook result. Various bases of solutions exist in vacuum, the particular bases of solutions we will focus on are the ‘elegant’ Gauss-Hermite beams, given by:
@@ V Propagation in homogeneous, isotropic media
V.0.1 ‘Elegant’ Hermite polynomial basis
@@
and it acts on $\exp\left(({\mathrm{i}}/{2})\bm{w}\cdot\bm{\Psi}_{w}\cdot\bm{w}\right)$. This ladder operator once again has the amplitude evolution built right into it. It is now clear that the complex Hermite polynomials are also given by both our ladder operators, and that the evolution of their profiles are identical for all three approaches.
We wish to point out that although the ‘elegant’ Gauss-Hermite beams are biorthogonal under the bilinear pairing inner product(section. B.3), this means that the elegant Gauss-Hermite beams do not neatly diagonalise in power, thus it is challenging to analyse the total power across the modes using this basis. An alternative basis of solutions that neatly diagonalises in power is the traditional Gauss-Hermite beam ansatz[pampaloni2004].
V.0.2 Traditional Hermite polynomial basis
@@
VI Relation to prior work by Pereversev
To the best our knowledge, prior work by Pereversev[Pere1998] is the most comprehensive and directly comparable attempt at beam tracing of an arbitrary profile. He does so by directly substituting in a Hermite polynomial basis of solutions into the Helmholtz equation, then analysing terms of various orders in the resulting expansion. This approach differs from our approach as it immediately prescribes a basis of solutions, whereas our results allow one to use any suitably chosen basis of solutions or even match boundary conditions exactly, simply by tweaking the ladder operators. An important feature to note is that our evolution equations for the beam envelope is decoupled from the profile evolution, whereas it is coupled to the profile evolution in Pereversev’s work. Whilst this may not have the biggest effect numerically, it certainly makes finding analytic solutions much more challenging relative to our approach.
@@ VI Relation to prior work by Pereversev
Pereversev does not treat eqs. 20, 21, 19, 24 separately, but rather considers them together by substituting in a specific basis of solutions. Upon substituting in the ansatz:
@@
$\begin{split}&\frac{\mathrm{d}\bm{\Psi}}{\mathrm{d}\tau}+\bm{\Psi}\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}(H)\bm{\Psi}+\bm{\Psi}\bm{\nabla}_{\bm{K}}\bm{\nabla}(H)+\bm{\nabla}\bm{\nabla}_{\bm{K}}(H)\bm{\Psi}\\ &+\bm{\nabla}\bm{\nabla}(H)=\bm{v}_{1}\bm{v}_{1}\cdot\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}(H)\cdot\bm{v}_{1}\bm{v}_{1}+\bm{v}_{2}\bm{v}_{2}\cdot\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}(H)\cdot\bm{v}_{2}\bm{v}_{2},\\ &\frac{\mathrm{dln}(A^{(0)})}{\mathrm{d}\tau}+\frac{\mathrm{d}\bm{K}}{\mathrm{d}\tau}\cdot\bm{\nabla}_{\bm{K}}\hat{e}_{m}\hat{e}^{*}_{m}+\hat{e}^{*}_{i}\frac{\partial}{\partial K_{\mu}}D_{im}\partial_{\mu}\hat{e}_{m}+\frac{1}{2}\Psi_{\mu\nu}\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}H\\ &+\mathrm{i}(2n+1)\bm{v}_{1}\cdot\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}(H)\cdot\bm{v}_{1}+\mathrm{i}(2m+1)\bm{v}_{2}\cdot\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}(H)\cdot\bm{v}_{2}=0.\end{split}$ (90)
@@
$\frac{\mathrm{d}^{n}}{\mathrm{d}\tau^{n}}(v_{1}^{\mu}\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}Hv_{2}^{\nu})|_{\tau=0}=0,$ (91)
implying that $v_{1}^{\mu}\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}Hv_{2}^{\nu}=0$ throughout the evolution if it is 0 at the boundary. Firstly, note that this solution is a simple product of Hermite polynomials without arguing that they diagonalise $\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H$, as if they did they would strictly be orthogonal to each other. However, we argued earlier that they have to be orthogonal in order for a simple product of Hermite polynomials to be the solution in our basis. However, it is still entirely possible that a sum of Hermite polynomial products could neatly reduce to a single product. Since Pereversev’s arguments in the Hermite polynomials differ from our arguments, it may still be the case that they are identical. In homogeneous media they are certainly identical.
@@
$v_{1}^{\mu}v_{2}^{\nu}\left(\frac{\mathrm{d}}{\mathrm{d}\tau}\left(\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}H\right)\right)_{w}=v_{1}^{\mu}v_{2}^{\nu}\left(T_{w,\mu\rho}\left(\partial K_{\rho}\frac{\partial}{\partial K_{\nu}}H\right)_{w}+T_{w,\nu\rho}\left(\partial K_{\rho}\frac{\partial}{\partial K_{\mu}}H\right)_{w}\right)$ (92)
@@
Since $\bm{v}_{I}$ ‘diagonalise’ $(\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}H)_{w}$ in vacuum, and remain orthogonal to each other throughout their evolution, Pereversev’s solution reproduces the exact same solutions that we obtained earlier in vacuum. This follows naturally from our derivation of Pereversev’s beam tracing equations from our profile evolution equation in eq. 90. Finding a situation in which our solutions do not coincide would prove that eq. 91 is not satisfied in general. We can only do so in a system where $\bm{v}_{I}$ do not diagonalise $(\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}H)_{w}$, and one good place to look for this is in propagation where the vectors rotate about $\hat{\bm{g}}$. Only by comparing the predictions of our theory with Pereversev’s in that context would one be able to investigate if our theories do in fact give identical results for the Gauss-Hermite boundary conditions. Although analytic work on obtaining the solutions in slab geometry[poli1999paraxial, Tsironis2006] has been done, the vectors do not rotate about $\hat{\bm{g}}$, nor is $(\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}H)_{w}$ non-diagonal in the first place. This means that the solutions they obtained will coincide with those obtained by this method, given the same boundary conditions. Unfortunately, we are unable to construct a physical situation where we are able to solve for analytically whilst rotating the vectors, thus this may be a task for numerical comparison.
@@
VII Implementation of theory
In this brief section, we discuss the numerical implementation of the theory developed in this paper. One would simply run traditional Gaussian beam tracing[HallChen2022], and then calculate the evolution of the ladder operators in post. Looking at eq. 70, we see that all we require are $\bm{v}_{I}(\tau)$ and $\bm{\Phi}(\tau)$. We can easily obtain $\bm{v}_{I}(\tau)$ from eq. 40 as it is a simple ODE. We will have to run a second beam tracing with the appropriately chosen $\bm{\Psi}_{w,\mathrm{tot}}=\bm{\Psi}_{w}+\bm{\Phi}$ in order to obtain $\bm{\Phi}(\tau)$, following the prescription given in eq. 47. We then substitute these into eq. 70 to obtain the beam profile at any point within the plasma, which should computationally be much more lightweight than solving the second-order PDE given in eq. 24.
For modelling dissipative effects, one has no choice but to solve eq. 72 in regions of dissipation, outside of which our exact solution would hold. Eq. 72 may be computationally simpler than solving the full eikonal, however it is still a complicated differential equation to solve in general. However, the ladder operator can be utilised in simple cases to obtain bases of solutions, for example in the case of linear width dependence, which may be the subject of further work.
VII.1 Further Work
Apart from theoretical work, work must go into computational implementation of our theory. The beam-tracing code of choice would Scotty[HallChen2022], as our theoretical model will simply reside as an additional module on top of its base beam-tracing capabilities for DBS. This work would hopefully enable Gauss-Hermite DBS analysis in the future, in order to provide more accurate signal analysis for imperfect antenna patterns.
In that same vein, we hope to utilise the results of this theory to develop a simplified theory of DBS, and maybe even cross-polarisation scattering(CPS), but with Gauss-Hermite modes – similar to the approach of previous work in DBS focused on the lowest order Gauss-Hermite beam[HallChen2022, HallChen2022b, HallChen2024]. We hope that this model will be computationally more efficient than evaluating the full volume integral of the reciprocity theorem, such that it is capable of providing real-time synthetic diagnostic analysis. Above all, we believe such a theory would give a clearer account of various factors that can affect the diagnostic signal, enabling accurate extraction of the turbulence spectrum from the signal, beyond just optimising diagnostic design. In order to do this, figuring out discontinuous boundary conditions would also be crucial.
VIII Conclusion
@@
We proceed to obtain a basis of solutions to the profile-evolution equation, for any choice of initial profile. These solutions were obtained using ladder operators, which can be used to set a wide range of polynomial boundary profiles. We focused on constructing ladder operators that generated functions which reduced to the Gauss-Hermite beams at the plasma boundary. Finally, we utilised our ladder operators to construct the exact solution to the propagation of a Gauss-Hermite beam through an inhomogeneous cold plasma using eq. 70. Most importantly however, this solution was derived directly from solving eq. 24 in full generality, thus we are confident in the accuracy of the result.
We finally demonstrated our ladder operators in action for propagation of a Gauss-Hermite beam through homogeneous media, before discussing future research direction, in particular, hopefully applying the results of this paper for improving real-time microwave diagnostics analysis.